OpenStudy (anonymous):

Maclaurin expansion series = x cos(x^2)

OpenStudy (anonymous):

if you don't already, you should half the maclaurin series for \(\cos(x)\) memorized

OpenStudy (anonymous):

easy to do since cosine is even it has all even exponents and it alternates

OpenStudy (anonymous):

then replace \(x\) by \(x^2\) in the previous maclaurin series finally, multiply your result by \(x\) i.e. kick up the exponent by 1 in each term

OpenStudy (anonymous):

if this is not clear, let me know, it is only 3 steps the first being a memory step

OpenStudy (anonymous):

so it would be x^2 cos x^3?

OpenStudy (anonymous):

oh no you are looking for the maclaurin series right? a power series

OpenStudy (anonymous):

do you know the maclaurin series for \(\cos(x)\) ?

OpenStudy (anonymous):

yeah and no i dont know

OpenStudy (anonymous):

ok you want to a) look it up b) derive it c) me to tell you what it is? either way which ever way, once you have it you should memorize it

OpenStudy (anonymous):

can you tell me what it is?

OpenStudy (anonymous):

sure and i will tell you how i memorize it too first, \(\cos(0)=1\) so it starts with 1 second, it alternates thirds, cosine is an even function, so all exponents are even we can write it out as \[1-\frac{x^2}{2!}+\frac{x^4}{4!}+\frac{x^6}{6!}...\]

OpenStudy (anonymous):

damn typo, it alternates!\[1-\frac{x^2}{2!}+\frac{x^4}{4!}-\frac{x^6}{6!}...\] that's better

OpenStudy (anonymous):

more succinctly written in sigma notation as \[\large \cos(x)=\sum_{n=0}^{\infty}\frac{(-1)^nx^{2n}}{(2n)!}\]

OpenStudy (anonymous):

memorize that one! also the one for \(\sin(x)\) and also \(e^x\)

OpenStudy (anonymous):

now for \(\cos(x^2)\) replace each \(x\) by \(x^2\) term by term it looks like \[1-\frac{x^4}{2!}+\frac{x^6}{4!}-\frac{x^8}{6!}...\]

OpenStudy (anonymous):

damn another typo square means double each exponent, let me try again

OpenStudy (anonymous):

\[1-\frac{x^4}{2!}+\frac{x^8}{4!}-\frac{x^{12}}{6!}...\] that's better

OpenStudy (anonymous):

more succinctly written in sigma notation as \[\cos(x^2)=\sum_{n=0}^{\infty}\frac{(-1)^nx^{4n}}{(2n)!}\]

OpenStudy (anonymous):

how we doing so far? one more step...

OpenStudy (anonymous):

So far so good. i understand everything you're saying. You break the problem down and everthing sound so clear.

OpenStudy (anonymous):

good then finally for \(x\cos(x^2)\) multiply all that mess by \(x\) term by term, which means really just add 1 to each exponent

OpenStudy (anonymous):

\[x(1-\frac{x^4}{2!}+\frac{x^8}{4!}-\frac{x^{12}}{6!}...\] \[x-\frac{x^5}{2!}+\frac{x^9}{4!}-\frac{x^{13}}{6!}...\]

OpenStudy (anonymous):

Thank you so much. I appreciate of what you did.

OpenStudy (anonymous):

in sigma notation i guess you can write this as \[\cos(x^2)=\sum_{n=0}^{\infty}\frac{(-1)^nx^{4n+1}}{(2n)!}\]

OpenStudy (anonymous):

oops i meant \(x\cos(x^2)\) but you know what i mean i hope yw